Company's existing costing system uses a single overhead rate. Based on total sales revenue from copy charges, to charge the cost of services department support activities to each size of machine. This service management has suggested that the copy charge should more accurately reflect the cost involved. The company accountants has decided to implement an activity based costing system and has obtained the following information about the support activities of the service department
A. Calculate the annual profit per machine for each of the three sizes of machine, using the current basis for charging the costs of support activities to machines
B. Calculate the annual profit per machine for each of the three sizes of machine using activity based costing
1
Expert's answer
2020-07-24T10:51:58-0400

A.

Let the income from all cars be $ 100,000 per year. and transaction volumes of 20,30,50 per machine, totaling 100 transactions. The price of one operation is $ 1000. Overhead costs 20 thousand dollars per year. Direct 30 thousand dollars a year. Then the waybills will be 20000/100 * 10 = 4000 for the first machine, 6000 for the second, 10,000 for the third. Similarly, direct costs 30,000 / 100 * 20 = 6000, 9000 for the second, 15000 for the third.


First profit:

1000 * 20-4000-6000 = 10,000

Profit for the second:

1000 * 30-6000-9000 = 30000-15000 = 15000

Profit for the third:

1000 * 50-10000-15000 = 50000-25000 = 25000


B.

Let the income from all cars be $ 100,000 per year. and transaction volumes of 20,30,50 per machine, totaling 100 transactions. The price of one operation is $ 1000. Overhead costs 20 thousand dollars per year. Direct 30 thousand dollars a year. Then the waybills will be for the first, second and third car 20,000/3 = 6,666.67. Similarly, direct costs for the first, second and third 30,000 / 3 = 10,000,


First profit:

1000 * 20-6 666.67-10 000 = 20 000-16 666.67 = 3333.33

Profit for the second:

1000 * 30-6 666.67-10 000 = 30 000-16 666.67 = 13 333.33

Profit for the third:

1000 * 50-6 666.67-10 000 = 50 000-16 666.67 = 33 333.33